OpenStudy (anonymous):

f(x) =√x + 6 and g(x) =1/ x^2-5 Find the function g ◦ f and determine its domain

OpenStudy (anonymous):

please help me solve

OpenStudy (lgbasallote):

g o f means g(f(x)) agree?

OpenStudy (anonymous):

ya..

OpenStudy (lgbasallote):

then is that \[f(x) = \sqrt x + 6\]or\[f(x) = \sqrt{x+6}\]

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

second one sorry